how many real solutions are there to the equation shown below?
x^2+2x+6=0

Found 2 solutions by lynnlo, ikleyn:
Answer by lynnlo(4176)

X=-2+2i√5/2

Answer by ikleyn(53419)

.
how many real solutions are there to the equation shown below?
x^2+2x+6=0
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


To answer this question, calculate the discriminant of the equation 

    d = b^2 - 4ac,    referring to the general form  ax^2 + bx + c = 0.


The discriminant is 

    d = 2^2 - 4*1*6 = 4 - 24 = -20.


The discriminant is negative - - - so, this equation has no real solutions.


ANSWER.  The number of real solutions is 0 (zero).

Solved.

It is how such questions should be treated.

----------------------------

The answer in the post by @lynnlo is incorrect,
so and therefore, ignore his post.